ABCD is a rhombus such that its its diagonals AC and BD intersect at M and satisfy BD=2AC. If the points D and M represent the complex no. 1+i and 2-i respectively, then which complex no. does A represent?

Grade:11

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

BD=2 AC

D=1+i

M=2-i

Using basic mathematics and the properties of rhombus we get the coordinates of A as 3-i/2 or 1-3i/2

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com

alley raphy
11 years ago

The points A,B,C and D represent the numbers Z1,Z2,Z3 and Z4.ABCD is rhombus;AC=2BD.if  Z2=2+i

,Z4=1-2i,find Z1 and Z3